All the dimensions of a certain rectangular solid are integers greater than 1. If the volume is 126 cubic inches and the height is 6 inches, what is the perimeter of the base?
(a) 10 (b)16 (c) 20 (d) 21 (e) 27

Answer by mananth(16946)

Volume = 126
height =6
area of base = 126/6 = 21 inches
L*W = 21
The dimensions of base = 7*3 =21
7" by 3"
The perimeter = 2*(L+W)
P= 2*(7+3)
P=20 inches